Kim traveled a total of 10 km. in 50 minutes.
Sasha traveled 10 km. in 55 minutes.
Barry did 10km. in 60 minutes.
What is their average speeds?

Answer by checkley75(3666)

10/(5/6)=60/5=12 kph.
10/(11/12)=120/11=10.9 kph.
10/1=10 kph.